OpenStudy (anonymous):

s^4-11s^2+28=0. find the real number solutions

OpenStudy (cwrw238):

factoring (s^2 - 7)(s^2 - 4) = 0 can you continue with this?

OpenStudy (anonymous):

So is would be s= to sqr rt of 7 and 2?

OpenStudy (cwrw238):

thats half right remember the negative square roots

OpenStudy (anonymous):

what about 164y^4=2

OpenStudy (cwrw238):

y^4 = 1/82 so what do you think?

OpenStudy (anonymous):

where did u get 82 from